# SePPO: Semi-Policy Preference Optimization for Diffusion Alignment

### v1

@@ 5.2.2 Analysis of Text-to-Video
We further validate SePPO on the text-to-video task in Table 3. Time-lapse video generation task is chosed and all the model are trained based on AnimateDiff. Our method achieves improvements across all metrics compared to both the vanilla and the fine-tuned AnimateDiff. We also visualize the video results in Figure 5, showing that SePPO achieves higher alignment with the text and better realism compared to the other methods.
@@ 5.2.3 Ablation Study
We perform an ablation study of our method in Table 4. When we remove AAF, meaning that all reference samples are treated as negative samples, the model’s performance drops significantly, demonstrating the effectiveness of AAF. Furthermore, when we replace the sign function ${\rm sgn}(x)$ in equation 9 with the indicator function $\mathds{1}(x>0)$, i.e., choosing to only learn from the winning image when the reference image has a higher probability of not being a losing image, and applying DPO when the reference image has a higher probability of being a losing image, we observe almost no change in performances. This proves that our method is able to filter out “insufficiently negative samples”–samples that are better than the current model’s distribution and subsequently boost performance. Furthermore, we experiment with different sampling strategies for the reference model with AAF. When we set the reference model’s sampling method to using either the initial checkpoint or the most recent saved checkpoint, we observe a performance drop in both cases. This indicates the effectiveness of our reference model sampling strategy.
@@
FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ FVD $\downarrow$
AnimateDiff 134.86 0.68 0.16 9.18 1608.41
SFT 129.14 0.65 0.17 9.25 1415.68
SePPO 115.32 0.61 0.20 9.36 1300.97
Table 4: Ablations on the Pick-a-Pic Validation Dataset.
Methods PickScore $\uparrow$ HPSv2 $\uparrow$ ImageReward $\uparrow$ Aesthetic $\uparrow$
w/o AAF 20.88 26.78 0.366 5.491
@@ 6.2 Conclusion
Without using reward models or human-annotated paired data, we have developed a Semi-Policy Preference Optimization (SePPO) method, which takes previous checkpoints as reference models and uses them to generate on-policy reference samples, which replace “losing images” in preference pairs. In addition, we design a strategy for reference model selection that expands the exploration in the policy space. Furthermore, instead of directly taking reference samples as negative examples, we propose an Anchor-based Adaptive Flipper to determine whether the reference samples are likely to be winning or losing images, which allows the model to selectively learn from the generated reference samples. In text-to-image benchmarks, SePPO achieved a $21.57$ PickScore, exceeding all previous approaches on the SD-1.5 model. In addition, SePPO performs better than SFT on text-to-video benchmarks.
Reproducibility Statement
For algorithms, we put the key parts (loss function) in Appendix B.2. We upload the main code for training to the supplementary material. The model checkpoints will be released shortly in https://github.com/DwanZhang-AI/SePPO. For datasets, we use open source datasets described in Section 5.1. For generated results, we upload generated videos to the supplementary material.
@@ Appendix A Theoretical Analysis
A.1 Proof of Theorem 4.1
@@
In the noise injection process, the variance of the images remains the same. ($\mathbf{x}_{0}$ and $\mathbf{x}_{t}$ have the same variance.) With the fact that the KL divergence between two Gaussian distributions with the identical variance is proportional to the Euclidean distance of their means, we have
@@
The prediction from model ${\theta_{1}}$ has a smaller KL distance compared to the prediction from model ${\theta_{2}}$. Thus, ${\theta_{1}}$ recovers better images and ${\theta_{1}}\succ{\theta_{2}}$ by the definition of performance measurement. As a result, ${\theta_{1}}$ has a higher probability of generating a good result $\mathbf{x}_{0}^{\theta_{1}}$.
@@ Appendix B Supplementary Experiments
B.1 Ablation Study on Iteration $K$
In this subsection, we perform an ablation study w.r.t. the number of iterations $K$. In Figure 6, we found that when changing the total number of iteration $K$ for saving checkpoints, relatively, the larger $K$ achieves better performance. However, the overall trend does not change significantly, which demonstrates the stability of SePPO on $K$.
Figure 6: Ablation study of $K=10$ and $K=20$.
B.2 Pseudocode
@@
[frame=shadowbox] def loss(target, model_pred, ref_pred, scale_term): #### START LOSS COMPUTATION #### if args.train_method == ’sft’: # SFT, casting for F.mse_loss loss = F.mse_loss(model_pred.float(), target.float(), reduction=”mean”) elif args.train_method == ’seppo’: # model_pred and ref_pred will be # (2 * LBS) x 4 x latent_spatial_dim x latent_spatial_dim # losses are both 2 * LBS # 1st half of tensors is preferred (y_w), # second half is unpreferred
@@
sign = torch.where(pos_losses_w ¿ 0, torch.tensor(1.0), torch.tensor(-1.0))
@@
scale_term = -0.5 * args.beta_dpo inside_term = scale_term * (model_diff - ref_diff) implicit_acc = (inside_term ¿ 0).sum().float() / inside_term.size(0) loss = -1 * (F.logsigmoid(inside_term)).mean()
@@
B.3 Visual Generation Examples
We present more text-to-visual generation results of SePPO and other methods. In Figure 7, we show the text-to-image generation results of SD-1.5, SPO, SPIN-Diffusion, and SePPO. In Figure 8 and Figure 9, we show the text-to-video generation results of AnimateDiff (Raw), SFT, and SePPO.
Figure 7: Text-to-image generation results of SD-1.5, SPO, SPIN-Diffusion, and SePPO. Prompts from left to right: (1) Pink Chihuahua in police suit; (2) Detailed Portrait Of A Disheveled Hippie Girl With Bright Gray Eyes By Anna Dittmann, Digital Painting, 120k, Ultra Hd, Hyper Detailed, Complimentary Colors, Wlop, Digital Painting; (3) Chic Fantasy Compositions, Ultra Detailed Artistic, Midnight Aura, Night Sky, Dreamy, Glowing, Glamour, Glimmer, Shadows, Oil On Canvas, Brush Strokes, Smooth, Ultra High Definition, 8k, Unreal Engine 5, Ultra Sharp Focus, Art By magali villeneuve, rossdraws, Intricate Artwork Masterpiece, Matte Painting Movie Poster; (4) winter owl black and white; (5) You are standing at the foot of a lush green hill that stretches up towards the sky. As you look up, you notice a beautiful house perched at the very top, surrounded by vibrant flowers and towering trees. The sun is shining brightly, casting a warm glow over the entire landscape. You can hear the sound of a nearby waterfall and the gentle rustling of leaves as a gentle breeze passes through the trees. The sky is a deep shade of blue, with a few fluffy clouds drifting lazily overhead. As you take in the breathtaking scenery, you can’t help but feel a sense of peace and serenity wash over you.
Figure 8: Text-to-video generation results of AnimateDiff (Raw), SFT, and SePPO. Prompt: “Time-lapse of night transitioning to dawn over a serene landscape with a reflective water body. It begins with a starry night sky and minimal light on the horizon, progressing through increasing light and a glowing horizon, culminating in a serene early morning with a bright sky, faint stars, and clear reflections in the water.
Figure 9: Text-to-video generation results of AnimateDiff (Raw), SFT, and SePPO. Prompt: “Time-lapse of aurora borealis over a night sky: starting with green arcs, intensifying with pronounced streaks, and evolving into swirling patterns. The aurora peaks in vivid hues before gradually fading into a homogeneous glow on a steadily brightened horizon.

### v2

@@ 5.2.3 Analysis of Text-to-Video
We further validate SSPO on the text-to-video task in Table 5. A time-lapse video generation task is chosen, and all the models are trained based on AnimateDiff. Our method achieves improvements across all metrics compared to the vanilla and the SFT AnimateDiff. We visualize video results in Figure 6, showing SSPO achieves higher alignment with the text and better realism compared to the others.
Figure 6: Text-to-video generation results of AnimateDiff (Raw), SFT, and SSPO. The prompt is: “Time-lapse of a lettuce seed germinating and growing into a mature plant. Initially, a seedling emerges from the soil, followed by leaves appearing and growing larger. The plant continues to develop…
@@ 5.2.4 Ablation Study
We perform an ablation study of modules in SSPO in Table 6. When we remove SSR, meaning that all the ERD samples are treated as negative samples, the model’s performance drops significantly, demonstrating the effectiveness of SSR. Furthermore, when we replace the sign function ${\rm sgn}(x)$ in equation 9 with the indicator function $\mathds{1}(x>0)$, i.e., choosing to only learn from the winning image when the ERD image has a higher probability of not being a losing image, and applying DPO when the ERD image has a higher probability of being a losing image, we observe almost no change in performances. This proves that our method is able to filter out “insufficiently negative samples”–samples that are better than the current model’s distribution and subsequently boost performance. Furthermore, we examined the necessity of RCR while retaining SSR. When we set the reference model’s sampling method to using either the initial checkpoint or the most recent saved checkpoint, we observe a performance drop in both cases. This indicates the effectiveness of our RCR strategy.
We then perform an ablation study w.r.t. the number of iterations $K$. In Figure 4, we found that when changing the total number of iterations $K$ for saving checkpoints, relatively, the larger $K$ achieves better performance. However, the overall trend does not change significantly, which demonstrates the stability of SSPO on $K$.
@@
FID $\downarrow$ PSNR $\uparrow$ FVD $\downarrow$ MTScore $\uparrow$ CHScore $\uparrow$
AnimateDiff 134.86 9.18 1608.41 0.402 65.37
SFT 129.14 9.25 1415.68 0.424 70.64
SSPO 115.32 9.36 1300.97 0.458 73.86
Table 6: Ablations on the Pick-a-Pic Validation Dataset. IR indicates Image Reward.
Methods PickScore $\uparrow$ HPSv2 $\uparrow$ IR $\uparrow$ Aesthetic $\uparrow$
w/o SSR 20.88 26.78 0.366 5.491
@@ 6.2 Conclusion
In this paper, we propose a Self-Sampling Preference Optimization (SSPO) method that integrates the advantages of SFT and DPO to effectively address the shortcomings of current pre-trained text-to-visual models in aligning with human requirements. We design a Random Checkpoint Replay strategy to construct effective paired data, and both experimental and theoretical analyses have demonstrated its superiority in mitigating overfitting risks and enhancing training stability. Additionally, the proposed Self-Sampling Regularization mechanism dynamically assesses the quality of generated responses to determine whether the samples are truly negative, thereby enabling flexible switching between DPO and SFT and preventing the model from learning biases from non-negative samples. Experimental results show that SSPO achieves significant performance improvements on both text-to-image and text-to-video tasks, validating its effectiveness in improving model generalization. In the future, we plan to further explore more refined negative sample generation strategies and evaluation mechanisms, to achieve even better performance in the broader field of AI-generated content.
Broader Impact Statement
This work contributes toward scalable and efficient post-training of generative models by reducing the dependence on expensive human annotations or external reward models. In particular, SSPO can democratize model alignment in domains where preference data is scarce, enabling broader adoption of diffusion-based generation techniques in resource-limited research or industry environments. However, the improved accessibility and controllability of powerful generative models also raises concerns. SSPO may inadvertently facilitate the misuse of aligned models in generating misleading or harmful content with higher perceptual quality and human alignment. Moreover, by enabling models to learn preference-aligned behaviors autonomously, there is a risk of encoding and amplifying biases present in the training data without human oversight.
@@ Appendix A Theoretical Analysis
A.1 Proof of Theorem 4.1 (Generalization)
We first define the generalization loss on the policy as $L(\pi_{\theta})\coloneqq\mathop{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}\mathcal{L}(\theta;\mathbf{x})$ and the empirical loss $\widehat{L}(\pi_{\theta})\coloneqq\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\theta;\mathbf{x}_{i})$, where $N$ is the data size. For the sake of conciseness, we omit the unnecessary superscripts and subscripts in $\mathbf{x}$. We then introduce the PAC-Bayesian theory Germain et al. (2016); Seeger (2002) as our Lemma A.1.
Lemma A.1
With a probability larger than $1-\Delta$, the following equation holds
$\mathop{\mathbb{E}}_{\pi_{\theta}\sim Q}\Big{[}{L}(\pi_{\theta})\Big{]}\leq\mathop{\mathbb{E}}_{\pi_{\theta}\sim Q}\Big{[}\widehat{{L}}(\pi_{\theta})\Big{]}+\sqrt{\frac{2D_{\rm KL}\big{(}Q||P\big{)}+\log\frac{4N}{\Delta^{2}}}{4N}},$ (11)
where $\Delta$ is the confidence radius, $P$ and $Q$ are prior and posterior distributions in the model space, respectively.
With the maximum entropy, Gaussian initialization is commonly used in policy learning Ahmed et al. (2019); Cen et al. (2022); Glorot & Bengio (2010). Thus, we make the following assumption that aligns well with practical implementations.
Assumption A.2
Assume the initial policies obey the Gaussian distribution $\mathcal{N}(0,\sigma_{0}^{2}\mathbf{I})$.
To simplify the notation in the proof, we shift the indices from $[0,K-1]$ to $[1,K]$, which means the $K$-th checkpoint is the latest one. In the $K$-th iteration, when the reference model is chosen as the last checkpoint, the model parameters are $\theta_{K}$. When the reference model is chosen uniformly distributed over all previous checkpoints, the expectation of the model parameters are $\frac{1}{K}\sum_{k=i}^{K}\pi_{\theta_{k}}$. The corresponding distributions are Dirac delta $\delta(\cdot)$ distributions. Thus, we denote the posterior distribution of SPIN $Q_{K}$ and SSPO $Q_{\rm UNI}$ as
$Q_{K}~{}={}$ $\delta(\pi_{\theta}-\pi_{\theta_{K}}),$ (12)
$Q_{\rm UNI}~{}={}$ $\frac{1}{K}\sum_{k=i}^{K}\delta(\pi_{\theta}-\pi_{\theta_{k}}).$
Using Lemma A.1, we have the upper bound of the generalization error of SPIN as follows
$\mathop{\mathbb{E}}_{\pi_{\theta}\sim Q_{K}}\Big{[}{L}(\pi_{\theta})\Big{]}\leq\widehat{{L}}(\pi_{\theta_{K}})+\sqrt{\frac{-2\log P(\pi_{\theta_{K}})+\log\frac{4N}{\Delta^{2}}}{4N}}\coloneqq\widehat{{L}}(\pi_{\theta_{K}})+C_{K}.$ (13)
The upper bound of the generalization error with uniform sampling is
$\mathop{\mathbb{E}}_{\pi_{\theta}\sim Q_{\rm UNI}}\Big{[}{L}(\pi_{\theta})\Big{]}\leq\frac{1}{K}\sum_{k=i}^{K}\widehat{{L}}(\pi_{\theta_{k}})+\sqrt{\frac{-\frac{2}{K}\sum_{k=1}^{K}\log P(\pi_{\theta_{k}})+\log\frac{4N}{\Delta^{2}}}{4N}}\coloneqq\frac{1}{K}\sum_{k=i}^{K}\widehat{{L}}(\pi_{\theta_{k}})+C_{\rm UNI}.$ (14)
We denote the difference $\epsilon$ between the upper bounds as follows
$\epsilon\coloneqq\widehat{{L}}(\pi_{\theta_{K}})-\frac{1}{K}\sum_{k=i}^{K}\widehat{{L}}(\pi_{\theta_{k}})+C_{K}-C_{\rm UNI}.$ (15)
When $K$ is large, the average loss could be arbitrarily close to the final loss. Thus, the difference term is dominated by the second part $\epsilon_{2}\coloneqq C_{K}-C_{\rm UNI}$.
Under Assumption A.2, the logarithmic prior distribution has a closed-form solution. We compare the difference between the KL divergences
$\epsilon_{\rm KL}\coloneqq-\log P(\pi_{\theta_{K}})+\frac{1}{K}\sum_{k=1}^{K}\log P(\pi_{\theta_{k}})=\frac{\|\pi_{\theta_{K}}\|^{2}}{2\sigma_{0}^{2}}-\frac{1}{K}\sum_{k=1}^{K}\frac{\|\pi_{\theta_{k}}\|^{2}}{2\sigma_{0}^{2}}.$ (16)
To calculate the norm of the policies, we trace back to its convergence performance. In SPIN Chen et al. (2024), it shows that $\pi_{\theta_{K}}$ exponentially convergences to $\pi_{\rm data}$ (the supervised label with one-hot distribution)
$\pi_{\theta_{K}}\propto\pi_{\theta_{K-1}}^{1-\frac{1}{\beta}}\pi_{\rm data}.$ (17)
Thus, with a large $K$, e.g., when it converges, $\|\pi_{\theta_{K}}\|^{2}=1$ achieves the upper bound of the second norm in the policy space. Let $\pi_{\theta}\in\mathbb{R}^{d}$, where $d$ is the dimension. Then, $\|\pi_{\theta_{0}}\|^{2}=\frac{1}{d}$. This can be easily achieved using the fact that a Gaussian distribution with standard deviation $\sigma$ has the same second norm as a uniform distribution with a radius of $\sigma\sqrt{d}$ in a high dimension. In general, it is a monotonic process. Without loss to generality, we simplify the convergence process as a linear process to show the result. Thus, we have
$\epsilon_{\rm KL}=\frac{1}{2\sigma_{0}^{2}}(1-\frac{1+\frac{1}{d}}{2})=\frac{1}{4\sigma_{0}^{2}}(1-\frac{1}{d})>0.$ (18)
Finally, we have
$\epsilon_{2}=\frac{\epsilon_{\rm KL}}{2N(C_{K}+C_{\rm UNI})}>0,$ (19)
which shows the uniformly sampling method has a lower confidence upper bound (The loss is smaller.) compared to the latest checkpoint selection method. Thus, the uniformly sampling method achieves better generalization ability with less overfitting risk.
A.2 Proof of Theorem 4.2
@@
In the noise injection process, the variance of the samples remains the same. ($\mathbf{x}_{0}$ and $\mathbf{x}_{t}$ have the same variance.) With the fact that the KL divergence between two Gaussian distributions with the identical variance is proportional to the Euclidean distance of their means, we have
@@
The prediction from model ${\theta_{1}}$ has a smaller KL distance compared to the prediction from model ${\theta_{2}}$. Thus, ${\theta_{1}}$ recovers better samples and ${\theta_{1}}\succ{\theta_{2}}$ by the definition of performance measurement. As a result, ${\theta_{1}}$ has a higher probability of generating a good result $\mathbf{x}_{0}^{\theta_{1}}$.
@@ Appendix B Supplementary Experiments
B.1 Results on the Parti-prompt Dataset.
From the results in Table.  7, it is evident that SSPO achieves the highest scores on key metrics such as PickScore and HPSv2, demonstrating its superior overall performance in meeting user preferences and generating visual content, even comparable to LS-Diff Zhang et al. (2024) which utilizes SDv2 as the base model. Meanwhile, SSPO maintains a high level of image quality (Aesthetic) and a stable image reward (IR), while still outperforming other methods in overall evaluations, which highlights its distinct advantage in balancing generation quality and alignment with human needs.
Table 7: Model Feedback Results on the Parti-prompt Dataset. IR indicates Image Reward. * indicates that the base model of LS-Diff is SDv2.
Methods PickScore $\uparrow$ HPSv2 $\uparrow$ IR $\uparrow$ Aesthetic $\uparrow$
SD-1.5 21.38 26.70 0.16 5.33
SFT 21.68 27.40 0.56 5.53
Diff-DPO 21.63 26.93 0.32 5.41
Diff-RPO 21.66 27.05 0.39 5.43
SPO 21.85 27.41 0.42 5.63
SPIN-Diff 21.91 27.58 0.51 5.78
LS-Diff - - 0.73 5.48
SSPO 21.93 27.79 0.58 5.64
B.2 Reward Distribution
(a) Distribution of SPIN and SPO
(b) Distribution of SFT and DPO
Figure 7: Reward distribution of several methods.
From the comparison of the reward distributions of different methods in the Figure. 7, it is evident that SSPO, compared to other methods (e.g., SPIN, SPO, SFT, DPO), demonstrates the following significant advantages:
Lower kurtosis in the reward distribution: As shown by the histogram and corresponding fitted curves, the SSPO distribution curve is flatter, and its peak is less pronounced than that of other methods. This indicates that SSPO’s reward distribution is more uniform during training, with fewer extremely high or low values, leading to better training stability.
Higher overall reward level: The bar chart illustrates that both the mean and median rewards obtained by SSPO are relatively higher, and the distribution is shifted to the right. This suggests that under the same training environment and task setup, SSPO can achieve better returns in most scenarios.
More balanced distribution: While other methods may produce extremely high or low rewards in certain cases, SSPO’s reward distribution is more concentrated. It not only attains a higher average reward but also effectively avoids excessive dispersion, thus enhancing overall robustness.
In summary, the reward distribution comparison in the figure indicates that SSPO, through more stable and efficient policy learning, achieves higher overall returns while maintaining a smoother distribution. It exhibits a distinctly comprehensive advantage across various metrics.
B.3 Human Evaluation
We conduct a human evaluation experiment based on the Pick-a-Pic validation dataset. Following D3PO (Yang et al., 2024), each image in the dataset was assessed by $5$ human raters. We then report the average percentage of images that received favorable evaluations in Figure. 8. The results demonstrate the superiority of SSPO compared to SPIN-diffusion.
Figure 8: Human evaluation result of SSPO and SPIN-diffusion
B.4 Step-wise Performance for SFT and SSPO
We then display the changes in PickScore during the training process of SSPO and SFT in Figure 9. SFT quickly converges and begins to fluctuate, while SSPO is able to steadily improve throughout the training process.
Figure 9: Step-aware performance for SFT and SSPO.
B.5 Pseudocode
@@
[frame=shadowbox] def loss(target, model_pred, ref_pred, scale_term): #### START LOSS COMPUTATION #### if args.train_method == ’sft’: # SFT, casting for F.mse_loss loss = F.mse_loss(model_pred.float(), target.float(), reduction="mean") elif args.train_method == ’SSPO’: # model_pred and ref_pred will be # (2 * LBS) x 4 x latent_spatial_dim x latent_spatial_dim # losses are both 2 * LBS # 1st half of tensors is preferred (y_w), # second half is unpreferred
@@
sign = torch.where(pos_losses_w > 0, torch.tensor(1.0), torch.tensor(-1.0))
@@
scale_term = -0.5 * args.beta_dpo inside_term = scale_term * (model_diff - ref_diff) implicit_acc = (inside_term > 0).sum().float() / inside_term.size(0) loss = -1 * (F.logsigmoid(inside_term)).mean()
@@
B.6 Visual Generation Examples
We present more text-to-visual generation results of SSPO and other methods. In Figure 10, we show the text-to-image generation results of SD-1.5, SPO, SPIN-Diffusion, and SSPO. In Figure 11 and Figure 12, we show the text-to-video generation results of AnimateDiff (Raw), SFT, and SSPO.
Figure 10: Text-to-image generation results of SD-1.5, SPO, SPIN-Diffusion, and SSPO. Prompts from left to right: (1) Pink Chihuahua in police suit; (2) Space Dog; (3) Chic Fantasy Compositions, Ultra Detailed Artistic, Midnight Aura, Night Sky, Dreamy, Glowing, Glamour, Glimmer, Shadows, Oil On Canvas, Brush Strokes, Smooth, Ultra High Definition, 8k, Unreal Engine 5, Ultra Sharp Focus, Art By magali villeneuve, rossdraws, Intricate Artwork Masterpiece, Matte Painting Movie Poster; (4) winter owl black and white; (5) You are standing at the foot of a lush green hill that stretches up towards the sky. As you look up, you notice a beautiful house perched at the very top, surrounded by vibrant flowers and towering trees. The sun is shining brightly, casting a warm glow over the entire landscape. You can hear the sound of a nearby waterfall and the gentle rustling of leaves as a gentle breeze passes through the trees. The sky is a deep shade of blue, with a few fluffy clouds drifting lazily overhead. As you take in the breathtaking scenery, you can’t help but feel a sense of peace and serenity wash over you.
Figure 11: Text-to-video generation results of AnimateDiff (Raw), SFT, and SSPO. Prompt: “Time-lapse of night transitioning to dawn over a serene landscape with a reflective water body. It begins with a starry night sky and minimal light on the horizon, progressing through increasing light and a glowing horizon, culminating in a serene early morning with a bright sky, faint stars, and clear reflections in the water.
Figure 12: Text-to-video generation results of AnimateDiff (Raw), SFT, and SSPO. Prompt: “Time-lapse of aurora borealis over a night sky: starting with green arcs, intensifying with pronounced streaks, and evolving into swirling patterns. The aurora peaks in vivid hues before gradually fading into a homogeneous glow on a steadily brightened horizon.